# Learning Heuristics for Transit Network Design and Improvement with Deep Reinforcement Learning

### v2

@@ Abstract
Transit agencies world-wide face tightening budgets. To maintain quality of service while cutting costs, efficient transit network design is essential. But planning a network of public transit routes is a challenging optimization problem. The most successful approaches to date use metaheuristic algorithms to search through the space of possible transit networks by applying low-level heuristics that randomly alter routes in a network. The design of these low-level heuristics has a major impact on the quality of the result. In this paper we use deep reinforcement learning with graph neural nets to learn low-level heuristics for an evolutionary algorithm, instead of designing them manually. These learned heuristics improve the algorithm’s results on benchmark synthetic cities with 70 nodes or more, and obtain state-of-the-art results when optimizing operating costs. They also improve upon a simulation of the real transit network in the city of Laval, Canada, by as much as 54% and 18% on two key metrics, and offer cost savings of up to 12% over the city’s existing transit network.
@@ Acknowledgements
We thank the Société de Transport de Laval for the map data and transportation data they provided, and the Agence Métropolitaine de Transport of Quebec for the origin-destination dataset they provided. This data was essential to the experiments presented in this paper. We also gratefully acknowledge the funding we received from NSERC in support of this work. And we thank Joshua Katz, the room-mate of the first author, for acting as a sounding board for ideas and giving feedback on our experimental designs.
@@ 1 INTRODUCTION
The NDP is the problem of designing a set of transit routes for a city that satisfy one or more objectives, such as meeting all travel demand and minimizing operating costs. It is an NP-complete problem that has commonalities with the travelling salesman problem (TSP) and vehicle routing problem (VRP), but is more complex than these famously-challenging problems. Since real-world cities typically have hundreds or even thousands of possible stop locations, analytical optimization approaches are infeasible. For this reason, the most successful approaches to the NDP to-date have been metaheuristic algorithms. Most such algorithms work by repeatedly applying one or more low-level heuristics that randomly modify a network, guiding this random walk towards better networks over many iterations by means of a metaheuristic such as natural selection (as in evolutionary algorithms) or metallic annealing (as in simulated annealing). Many different metaheuristics for guiding the search have been proposed, as have a range of low-level heuristics for transit network design specifically. While these algorithms can find good solutions in many cases, real-world transit networks are still commonly designed by hand (Durán-Micco and Vansteenwegen, 2022).
@@
In our prior work (Holliday and Dudek, 2023), we trained a graph neural net (GNN) to plan a transit network from scratch, and showed that using this to generate an initial transit network improved the quality of networks found by two metaheuristic algorithms. That policy assembled transit routes by concatenating shortest paths in the city’s street network that share end-points, using the neural net to select the next shortest path to be included at each step. By repeating this process for each route that is needed, it assembled a complete transit network.
@@
In the present work, we expand on those results in several ways. We present new results obtained with an improved GNN model and an improved evolutionary algorithm. We also perform ablation studies to analyze the importance of different components of our neural-evolutionary hybrid algorithm. As part of these ablations, we consider a novel unlearned low-level heuristic, where shortest paths with common end-points are joined uniformly at random instead of according to the GNN’s output. Interestingly, we find that in the narrow case where we are only concerned with minimizing passenger travel time, this heuristic outperforms both the baseline heuristics of Nikolić and Teodorović (2013) and our GNN heuristic; in most other cases, however, our GNN heuristic performs better.
We compare the performance of our learned and unlearned heuristics with other results on the widely-used Mandl (Mandl, 1980) and Mumford (Mumford, 2013a) benchmark cities. We find that the unlearned heuristic obtains results comparable to the state of the art when minimizing only the passengers’ travel time, while the GNN heuristic outperforms the previous state of the art by as much as 13% when minimizing only the cost to operators.
@@ 2.1 Deep Learning for Optimization Problems
Deep learning refers to machine learning techniques that involve “deep” artificial neural nets - that is, neural nets with many hidden layers. Deep reinforcement learning refers to the use of deep neural nets in reinforcement learning (RL). RL is a branch of machine learning in which a learning system, such as a neural net, chooses actions and receives a numerical “reward” for each action, and learns to choose actions that maximize total reward. As documented by  Bengio et al. (2021), there is growing interest in the application of deep learning, and deep RL specifically, to combinatorial optimization (CO) problems such as the TSP.
Vinyals et al. (2015) proposed a deep neural net model called a Pointer Network, and trained it via supervised learning to solve TSP instances. Subsequent work, such as that of Dai et al. (2017), Kool et al. (2019), and Sykora et al. (2020), has built on this. These works use similar neural net models, in combination with reinforcement learning algorithms, to train neural nets to construct CO solutions. These have attained impressive performance on the TSP, the VRP, and other CO problems. More recently, Mundhenk et al. (2021) train an recurrent neural net (RNN) via RL to construct a starting population of solutions to a genetic algorithm, the outputs of which are used to further train the RNN. Fu et al. (2021) train a model on small TSP instances and present an algorithm that applies the model to much larger instances. Choo et al. (2022) present a hybrid algorithm of Monte Carlo Tree Search and Beam Search that draws better sample solutions for the TSP and Capacitated (CVRP) from a neural net policy like that of Kool et al. (2019).
@@
Some work has considered training neural nets to choose the search moves to be made at each step of an improvement method (Hottung and Tierney, 2019; Chen and Tian, 2019; d O Costa et al., 2020; Wu et al., 2021; Ma et al., 2021), and Kim et al. (2021) train one neural net to construct a set of initial solutions, and another to modify and improve them. This work has shown impressive performance on the TSP and other CO problems. Our work belongs to this family, in that we train a neural net to solve instances of the transit network design problem, and then use it to modify solutions within an improvement method.
In most of the above work, the neural net models used are graph neural nets, a type of neural net model that is designed to operate on graph-structured data (Bruna et al., 2013; Kipf and Welling, 2016; Defferrard et al., 2016; Duvenaud et al., 2015). These have been applied in many other domains, including the analysis of large web graphs (Ying et al., 2018), the design of printed circuit boards (Mirhoseini et al., 2021), and the prediction of chemical properties of molecules (Duvenaud et al., 2015; Gilmer et al., 2017). An overview of GNNs is provided by Battaglia et al. (2018). Like many CO problems, the transit network design problem lends itself to being described as a graph problem, so we use GNN models here as well.
In most CO problems, it is difficult to find globally optimal solutions but easier gauge the quality of a given solution. As  Bengio et al. (2021) note, this makes reinforcement learning a natural fit to CO problems. Most of the work cited in this section uses RL methods to train neural net models. We do the same here, applying a method similar to that of Kool et al. (2019) to the NDP.
@@ 2.2 Optimization of Public Transit
The NDP is an NP-complete problem (Quak, 2003), meaning that it is impractical to find optimal solutions for most cases. While analytical optimization and mathematical programming methods have been successful on small instances (van Nes, 2003; Guan et al., 2006), they struggle to realistically represent the problem (Guihaire and Hao, 2008; Kepaptsoglou and Karlaftis, 2009). Metaheuristic approaches, as defined by Sörensen et al. (2018), have thus been more widely applied.
@@ 3 TRANSIT NETWORK DESIGN PROBLEM
    $\mathcal{R}$ must satsify all demand, providing some path over transit between every pair of nodes $(i,j)\in\mathcal{N}$ for which $D_{ij}>0$.
@@ 3.1 Markov Decision Process Formulation
The starting state is $s_{0}=(\mathcal{R}_{0}=\{\},r_{0}=[])$. At a high level, the MDP alternates at every timestep between two modes: on odd-numbered timesteps, the agent selects an extension to the route $r_{t}$ that it is currently planning; on even-numbered timesteps, the agent chooses whether or not to stop extending $r_{t}$ and add it to the set of finished routes.
@@
The episode ends when the $S$-th route is added to $\mathcal{R}$: that is, if $|\mathcal{R}_{t+1}|=S$, the episode ends at timestep $t$. The output transit network is set to $\mathcal{R}=\mathcal{R}_{t+1}$, and the final reward is $R_{t_{\text{end}}}=-C(\mathcal{C},\mathcal{R})$. At all prior steps, $R_{t}=0$.
@@
3.2 Cost Function
Our NDP cost function has three components. The cost to passengers, $C_{p}$, is the average passenger trip time over the network:
Where $\tau_{\mathcal{R}ij}$ is the time of the shortest transit trip from $i$ to $j$ given $\mathcal{R}$, including a time penalty $p_{T}$ for each transfer the rider must make between routes.
$C_{o}(\mathcal{C},\mathcal{R})=\sum_{r\in\mathcal{R}}\tau_{r}$ (7)
To enforce the constraints on $\mathcal{R}$, we use a third term $C_{c}$, which is the fraction of node pairs $(i,j)$ with $D_{ij}>0$ that are not connected by $\mathcal{R}$, plus a measure of by how many stops $|r|>MAX$ or $|r|<MIN$ across all routes. The cost function is then:
@@ 4.1 Learning to Construct a Network
Figure 2: A schematic of our proposed route-planning method, outlining the route-construction MDP and showing the role of the policy network’s three components: the GAT backbone, $\textup{NN}_{ext}$, and $\textup{NN}_{halt}$.
In order to learn heuristics for transit network design, we first train a GNN policy $\pi_{\theta}(a|s)$, parameterized by $\theta$, to maximize the cumulative return $G_{t}$ on the construction MDP described in subsection 3.1. By then following this learned policy on the MDP for some city $\mathcal{C}$, stochastically sampling $a_{t}$ at each step with probabilities given by $\pi(\cdot|s_{t})$, we can obtain a transit network $\mathcal{R}$ for that city. We refer to this construction algorithm using the learned policy as “learned construction”. Because learned construction stochastically samples actions, we can run it multiple times to generate multiple networks for a city, and then pick the lowest-cost one. We will often do this by running learned construction 100 times and choosing the best network; we denote this procedure “LC-100”.
@@
The GAT outputs node embeddings $\mathbf{y}_{i}$, which are operated on by one of two “head” neural nets, depending on the timestep: $\textup{NN}_{ext}$ for choosing among extensions when the timestep $t$ is odd, and $\textup{NN}_{halt}$ for deciding whether to halt when $t$ is even. Figure 2 is a schematic illustration of the function of each of these three components in the route-planning MDP. The full details of the neural architecture are presented in Appendix A. We have released the code for our method and experiments to the public [^2] Available at https://www.cim.mcgill.ca/~mrl/hgrepo/transit_learning/.

### v3

@@ Abstract
Transit agencies world-wide face tightening budgets. To maintain quality of service while cutting costs, efficient transit network design is essential. But planning a network of public transit routes is a challenging optimization problem. The most successful approaches to date use metaheuristic algorithms to search through the space of possible transit networks by applying low-level heuristics that randomly alter routes in a network. The design of these low-level heuristics has a major impact on the quality of the result. In this paper we use deep reinforcement learning with graph neural nets to learn low-level heuristics for an evolutionary algorithm, instead of designing them manually. These learned heuristics improve the algorithm’s results on benchmark synthetic cities with 70 nodes or more, and obtain state-of-the-art results when optimizing operating costs. They also improve upon a simulation of the real transit network in the city of Laval, Canada, by as much as 52% and 24% on two key metrics, and offer cost savings of up to 17% over the city’s existing transit network.
@@ Acknowledgements
We thank the Société de Transport de Laval for the map data and transportation data they provided, and the Agence Métropolitaine de Transport of Quebec for the origin-destination dataset they provided. This data was essential to the experiments presented in this paper. We also gratefully acknowledge the funding we received from NSERC in support of this work. And we thank Joshua Katz for acting as a sounding board for ideas and giving feedback on our experimental designs.
@@ 1 INTRODUCTION
The NDP is the problem of designing a set of transit routes for a city that satisfy one or more objectives, such as meeting all travel demand and minimizing operating costs. It is an NP-complete problem (Quak, 2003), and since real-world cities typically have hundreds or even thousands of possible stop locations, analytical optimization approaches are usually infeasible. The most successful approaches to the NDP to-date have been metaheuristic algorithms. Most such algorithms work by repeatedly applying one or more low-level heuristics that randomly modify a network, guiding this random walk towards better networks over many iterations by means of a metaheuristic such as natural selection (as in evolutionary algorithms) or metallic annealing (as in simulated annealing). Many different metaheuristics for guiding the search have been proposed, as have a range of low-level heuristics for transit network design specifically. While these algorithms can find good solutions in many cases, real-world transit networks are still commonly designed by hand (Durán-Micco and Vansteenwegen, 2022).
@@
In our prior work (Holliday and Dudek, 2023), we trained a graph neural net (GNN) to plan a transit network from scratch, and showed that using this to generate an initial transit network improved the quality of networks found by two metaheuristic algorithms. That policy assembled transit routes by concatenating shortest paths in the city’s street network that share end-points, using the neural net to select the next shortest path to be included at each step. By repeating this process for each route that is needed, it assembled a complete transit network. Key to the success of this approach was our novel algorithm for computing a heuristic over the shortest paths at each construction step from the neural net’s outputs.
@@
In the present work, we expand on those results in several ways. We present new results obtained with an improved GNN model trained via Proximal Policy Optimization (PPO) instead of the older REINFORCE algorithm, as well as using an improved evolutionary algorithm. We also perform ablation studies to analyze the importance of different components of our neural-evolutionary hybrid algorithm. As part of these ablations, we consider a novel unlearned low-level heuristic, where shortest paths with common end-points are joined uniformly at random instead of according to the GNN’s output. Interestingly, we find that in the narrow case where we are only concerned with minimizing passenger travel time, this heuristic outperforms both the baseline heuristics of Nikolić and Teodorović (2013) and our GNN heuristic; in most other cases, however, our GNN heuristic performs better.
We compare the performance of our learned and unlearned heuristics with other results on the widely-used Mandl (Mandl, 1980) and Mumford (Mumford, 2013a) benchmark cities. We find that the unlearned heuristic obtains results comparable to the state of the art when minimizing only the passengers’ travel time, while the GNN heuristic outperforms the previous state of the art by as much as 3% when minimizing only the cost to operators.
@@ 2.1 Deep Learning for Optimization Problems
Deep learning refers to machine learning techniques that involve “deep” artificial neural nets - that is, neural nets with many hidden layers. Deep reinforcement learning refers to the use of deep neural nets in reinforcement learning (RL). RL is a branch of machine learning in which a learning system, such as a neural net, chooses actions and receives a numerical “reward” for each action, and learns to choose actions that maximize total reward. As documented by  Bengio et al. (2021), there is growing interest in the application of deep learning, and deep RL specifically, to combinatorial optimization (CO) problems.
Vinyals et al. (2015) proposed a deep neural net model called a Pointer Network, and trained it via supervised learning to solve instances of the travelling salesman problem (TSP). Subsequent work, such as that of Dai et al. (2017), Kool et al. (2019), and Sykora et al. (2020), has built on this. These works use similar neural net models, in combination with reinforcement learning algorithms, to train neural nets to construct CO solutions. These have attained impressive performance on the TSP, the vehicle routing problem (VRP), and other CO problems. More recently, Mundhenk et al. (2021) train an recurrent neural net (RNN) via RL to construct a starting population of solutions to a genetic algorithm, the outputs of which are used to further train the RNN. Fu et al. (2021) train a model on small TSP instances and present an algorithm that applies the model to much larger instances. Choo et al. (2022) present a hybrid algorithm of Monte Carlo Tree Search and Beam Search that draws better sample solutions for the TSP and Capacitated (CVRP) from a neural net policy like that of Kool et al. (2019).
@@
Some work (Hottung and Tierney, 2019; Chen and Tian, 2019; da Costa et al., 2020; Wu et al., 2021; Ma et al., 2021), has considered training neural nets to choose the search moves to be made at each step of an improvement method. meanwhile Kim et al. (2021) train one neural net to construct a set of initial solutions, and another to modify and improve them. And more recently, Ye et al. (2024) train a neural net to provide a heuristic score for choices in CO problems in the context of an ant colony optimization algorithm. This work has shown impressive performance on the TSP, VRP, and similar CO problems.
In most of the above work, the neural net models used are graph neural nets, a type of neural net model that is designed to operate on graph-structured data (Bruna et al., 2013; Kipf and Welling, 2016; Defferrard et al., 2016; Duvenaud et al., 2015). These have been applied in many other domains, including the analysis of large web graphs (Ying et al., 2018), the design of printed circuit boards (Mirhoseini et al., 2021), and the prediction of chemical properties of molecules (Duvenaud et al., 2015; Gilmer et al., 2017). An overview of GNNs is provided by Battaglia et al. (2018). Like many CO problems, the NDP lends itself to being described as a graph problem, so we use GNN models here as well.
CO problems also have in common that it is difficult to find a globally optimal solution but easier to gauge the quality of a given solution. As  Bengio et al. (2021) note, this makes reinforcement learning a natural fit to CO problems. Most of the work cited in this section uses RL methods to train neural net models.
@@ 2.2 Optimization of Public Transit
The above work all concerns a small set of classical CO problems including the TSP and VRP. Like the NDP, some of these are NP-complete, and all can be described as graph problems. But the NDP resembles these classical problems much less than they do each other: transit routes can interact by transferring passengers, and the space of solutions for a problem instance of given size is much vaster in the NDP. Take as an example the small Mumford0 benchmark city described in Table 1. With 30 nodes there are $30!\approx 2.7\times 10^{32}$ possible TSP tours, but for 12 routes with at least 2 and at most 15 stops each, there are approximately $2.3\times 10^{11}$ possible routes and so (${2.3\times 10^{11}\choose 12}\approx 4.3\times 10^{127}$ possible transit networks, a factor of $10^{95}$ more.
The difference is further shown by the fact that the state-of-the-art on classic problems like the TSP uses analytical and mathematical programming methods (such as Applegate et al. (2001)), but on the NDP such methods are not used. The NDP’s differences necessitate a special treatment, and so all recent published work on the NDP deals with the NDP alone (Mumford, 2013a; John et al., 2014; Kılıç and Gök, 2014; Islam et al., 2019; Ahmed et al., 2019; Lin and Tang, 2022; Hüsselmann et al., 2023), and we do the same here.
While analytical and mathematical programming methods have been successful on small instances (van Nes, 2003; Guan et al., 2006), they struggle to realistically represent the problem (Guihaire and Hao, 2008; Kepaptsoglou and Karlaftis, 2009). Metaheuristic approaches, as defined by Sörensen et al. (2018), have thus been more widely applied.
@@ 3 TRANSIT NETWORK DESIGN PROBLEM
    $\mathcal{R}$ must satisfy all demand, providing some path over transit between every pair of nodes $(i,j)\in\mathcal{N}$ for which $D_{ij}>0$.
@@
3.1 Cost Function
We take the NDP cost function to have three components. The cost to passengers, $C_{p}$, is the average passenger trip time over the network:
$C_{p}(\mathcal{C},\mathcal{R})=\frac{\sum_{i,j}D_{ij}(\delta_{\mathcal{R}ij}\tau_{\mathcal{R}ij}+(1-\delta_{\mathcal{R}ij})2\max_{kl}T_{kl})}{\sum_{i,j}D_{ij}}$ (2)
Where $\tau_{\mathcal{R}ij}$ is the time of the shortest transit trip from $i$ to $j$ given $\mathcal{R}$, including a time penalty $p_{T}$ for each transfer the rider must make between routes, and $\delta_{\mathcal{R}ij}$ is a delta function that takes value 1 if $\mathcal{R}$ provides a path from $i$ to $j$, and 0 if not. If $\mathcal{R}$ satisfies constraint 1, it connects all pairs of nodes, so for all feasible networks the above equation simplifies to:
Otherwise, each demand that is unsatisfied is treated as taking twice the largest inter-node driving distance to be satisfied.
$C_{o}(\mathcal{C},\mathcal{R})=\sum_{r\in\mathcal{R}}\tau_{r}$ (4)
To enforce the constraints on $\mathcal{R}$, we use a third term $C_{c}$, itself the sum of three terms:
$C_{c}(\mathcal{C},\mathcal{R})=F_{un}+F_{s}+0.1\delta_{v}$ (5)
Here, $F_{un}$ is the fraction of node-pairs $(i,j)$ with $D_{ij}>0$ for which $\mathcal{R}$ provides no path: a measure of how many possible violations of constraint 1 really occur in $\mathcal{R}$. $F_{s}$ is a similar measure for constraint 3, but it is proportional to the actual number of stops more than $MAX$ or less than $MIN$ that each route has. Specifically:
$F_{s}=\frac{\sum_{r\in\mathcal{R}}\max(0,MIN-|r|,|r|-MAX)}{S*MAX}$ (6)
$\delta_{v}$ is a delta function that takes value 0 if $F_{un}=0$ and $F_{s}=0$, and takes value 1 otherwise.
We use fractional measures $F_{un}$ and $F_{s}$ instead of absolute measures so that $C_{c}$ will tend to fall in the range $[0,1]$ regardless of the size of the city graph. This is desirable for reasons relating to the numerical stability of neural net training. However, it has the drawback that if the city graph $C$ has very many nodes (large $n$), and if $\mathcal{R}$ violates only a few constraints, $F_{un}$ and $F_{s}$ may become vanishingly small compared to $C_{o}$ and $C_{p}$. This could lead to algorithms ignoring small numbers of constraint violations. To prevent this, we include the term $0.1\delta_{v}$, which ensures that if any constraints are violated, $C_{c}$ cannot be less than 0.1, but it will be 0 if all constraints are respected. This ensures that even one violation will have a significant impact on overall cost $C(\mathcal{C},\mathcal{R})$, no matter the size of the city graph.
$C_{c}$ does not penalize violations of constraints 2 and 4, because these constraints are enforced directly by the design of the Markov Decision Process (MDP) detailed in subsection 3.2.
The cost function is then:
@@ 3.2 Markov Decision Process Formulation
The starting state is $s_{1}=(\mathcal{R}_{1}=\{\},r_{1}=[])$. At a high level, the MDP alternates at every timestep between two modes: on odd-numbered timesteps, the agent selects an extension to the route $r_{t}$ that it is currently planning; on even-numbered timesteps, the agent chooses whether or not to stop extending $r_{t}$ and add it to the set of finished routes.
@@
The episode ends when the $S$-th route is added to $\mathcal{R}$: that is, if $|\mathcal{R}_{t+1}|=S$, the episode ends at timestep $t$. The output transit network is set to $\mathcal{R}=\mathcal{R}_{t+1}$. The reward function is defined as the decrease in cost between each consecutive pair of timesteps:
$R_{t}=C^{\prime}(\mathcal{C},\mathcal{R}_{t}\cup\{r_{t}\})-C^{\prime}(\mathcal{C},\mathcal{R}_{t+1}\cup\{r_{t+1}\})$ (12)
Note that this means that on even timesteps (when halt-or-continue actions are taken), $R_{t}=0$ because $\mathcal{R}_{t}$ and $r_{t}$ do not change as a result of a halt-or-continue action.
In the above equation, $C^{\prime}((\mathcal{C},\mathcal{R})=C(\mathcal{C},\mathcal{R})-\beta F_{s}$; that is, the cost function without the $F_{s}$ component that penalizes violations of constraint 3. We remove this component because, by construction of the MDP, the networks it produces always satisfy constraint 3, so there is no need to penalize violations of it that may occur before it is complete.
Figure 2: A flowchart of the transit network construction process defined by our MDP. Blue boxes indicate points where the timestep $t$ is incremented and the agent selects an action. Red nodes are the beginning and ending of the process. Green nodes are hard-coded decision points. Orange nodes show updates to the state $\mathcal{S}$ and action space $\mathcal{A}$.
@@ 4.1 Learning to Construct a Network
In order to learn heuristics for transit network design, we first train a GNN policy $\pi_{\theta}(a|s)$, parameterized by $\theta$, to maximize the cumulative return $G_{t}$ on the construction MDP described in subsection 3.2. By then following this learned policy on the MDP for some city $\mathcal{C}$, stochastically sampling $a_{t}$ at each step with probabilities given by $\pi(\cdot|s_{t})$, we can obtain a transit network $\mathcal{R}$ for that city. We refer to this construction algorithm using the learned policy as “learned construction”. Because learned construction samples actions stochastically, we can run it multiple times to generate multiple networks for a city, and then pick the lowest-cost one. We will often do this by running learned construction 100 times and choosing the best network; we denote this procedure “LC-100”. As noted in subsection 2.2, the NDP has a much vaster space of possible solutions than problems like the TSP or VRP. To make this vast problem tractable, we found it necessary to “factorize” the computation of heuristic scores for each shortest-path extension to a route. Our policy net thus outputs a scalar score $o_{ij}$ for each node pair $(i,j)$ in the graph at each step. From these, a score for each extension $o_{a}$ is computed by summing over $o_{ij}$ for all $(i,j)$ that would become directly connected by appending path $a$ to the current route $r_{t}$:
$o_{a}=\sum_{i\in r_{t}}\sum_{j\in a}o_{ij}+\sum_{k\in a}\sum_{l\in a}o_{kl}$
Where $o_{ii}=0$ for all $i$. This factorization makes the learning objective “estimate the benefit of directly connecting nodes $i$ and $j$ with a given travel time”, rather than the much more complex “estimate the quality of adding path $a$ to $r_{t}$”. This objective is easier to learn, and reduces the needed complexity of the neural net architecture.
@@
The GAT outputs node embeddings $\mathbf{y}_{i}$, which are operated on by one of two “head” neural nets, depending on the timestep: $\textup{NN}_{ext}$ and $\textup{NN}_{halt}$. $\textup{NN}_{ext}$ computes the $o_{ij}$ node-pair scores from the node embeddings $\mathbf{y}_{i}$ and the driving time from $i$ to $j$ along $r_{t}|a$ when the timestep $t$ is odd. $\textup{NN}_{halt}$ decides whether to halt when $t$ is even, based on $\mathbf{y}_{i}$ and on other statistics of the city $\mathcal{C}$ and partial network $\mathcal{R}_{t},r_{t}$. The full architectural details of these components are presented in Appendix A, along with the details of the input node and edge features.
We have released the code for our method and experiments to the public222 Available at https://github.com/AHolliday/transit_learning.